### v1

# P2LHAP:Wearable sensor-based human activity recognition, segmentation and forecast through Patch-to-Label Seq2Seq Transformer

Abstract
Traditional deep learning methods struggle to simultaneously segment, recognize, and forecast human activities from sensor data. This limits their usefulness in many fields such as healthcare and assisted living, where real-time understanding of ongoing and upcoming activities is crucial. This paper introduces P2LHAP, a novel Patch-to-Label Seq2Seq framework that tackles all three tasks in a efficient single-task model. P2LHAP divides sensor data streams into a sequence of "patches", served as input tokens, and outputs a sequence of patch-level activity labels including the predicted future activities. A unique smoothing technique based on surrounding patch labels, is proposed to identify activity boundaries accurately. Additionally, P2LHAP learns patch-level representation by sensor signal channel-independent Transformer encoders and decoders. All channels share embedding and Transformer weights across all sequences. Evaluated on three public datasets, P2LHAP significantly outperforms the state-of-the-art in all three tasks, demonstrating its effectiveness and potential for real-world applications.
Index Terms:
sensor-based human activity recognition, activity segmentation, activity forecast, deep learning, Patch-to-Label Seq2Seq framework.
I Introduction
The intersection of deep learning and wearable sensor technology for real-time human activity recognition has emerged as a prominent research focus in recent years. This area holds significant promise for applications in medical health, assisted living, smart homes, and other domains. In various practical scenarios, there is a need for perceptual methods to delineate activity boundaries, identify activity categories, and predict forthcoming activities from the continuous stream of sensor data. As a result, human activity perception encompasses three pivotal tasks: human activity segmentation (HAS)[51], activity recognition (HAR)[50], and activity forecast (HAF)[49]. While extensive research[42, 43] has addressed each of these tasks individually, limited attention has been devoted to simultaneously addressing all three tasks in a cohesive framework.
Several approaches exist for concurrently achieving activity segmentation, identification, and forecast, as illustrated in Fig. 1. The first approach Fig. 1(a) employs three independent single-task learning models in parallel. These models output the category of the current activity, the boundary of the current activity, and the category of the next activity. However, this method has drawbacks, including disparate data annotation requirements for different models, the need for additional policy compatibility model outputs, and a lack of information sharing between models, leading to inefficient use of computing resources.
A second approach involves multi-task learning in Fig. 1(b), where multiple interrelated tasks share data and encoders, learning a common data representation. Subsequently, multiple lightweight single-task models [5, 6, 25] are employed. While this approach often enhances multitasking effectiveness, it introduces increased model complexity compared to single-task models. Managing this complexity may necessitate additional computational resources and careful architectural design. Furthermore, these multitask models can be more sensitive to hyperparameter choices, demanding meticulous tuning to achieve optimal performance across all tasks.
Different from the first two methods, as shown in Fig. 1(c), this paper proposes a novel Patch-to-Label Seq2Seq framework that tackles all the three human activity perception tasks in a single-task model, named P2LHAP. P2LHAP adopts the encoder-decoder framework, the input is a patch sequence of sensor data, and the output is a sequence of patch-level activity label, where a patch is a short segment of sensor data points. With patch-level labels, as shown in Fig. 1(c), it can simply Merge adjacent identical labels to determine the category of the active boundary. In addition, using the encoder-decoder architecture of seq2seq, It is easy to output future activity token sequences, making it simple to predict activity label. The architecture handling capability of variable-length sensor activity sequences is excellent, and it exhibits high adaptability to sequences of indefinite length. Moreover, it comprehensively considers all information within the input sensor signal sequence, effectively retaining the contextual information of the sensor data stream.
Using patches offers several advantages over individual data points and sliding windows. Compared to using single data points[16, 6], using patches can reduce the impact of noise, allowing feature extraction within local intervals[41]. Most existing activity recognition research uses one or more fixed-length slides Windows segment the data stream, assuming each window contains a complete activity[5]. However, due to the changing duration of the activity, the window length is very difficult to set. If it is too small, the window does not contain enough information; if it is too large, a window contains activities of multiple different categories, both of which will reduce the recognition accuracy [53]. The patching strategy can circumvent this problem well, because we assume that an activity can contain any number of patches, and encoder-decoder architectures such as tranfomer are good at handling variable-length sequences, making it feasible to identify activities of arbitrary length.
Figure 1: There are different architectures for multitasking approaches as follows: (a) The data is partitioned using sliding windows, followed by separate training of the respective models. These trained models are subsequently utilized for recognition, segmentation, and forecast. (b) Multi-task methods learn models for multiple tasks through a shared encoder and generate separate results for each task. (c) We propose a multi-task P2L Seq2Seq model specifically designed to generate patch-level active label sequences.
We introduce a channel-independent transfomer as the encoder-decoder of P2LHAP. Although transformers have achieved great results in the fields of NLP[7] and CV[55], recent research shows that classic transformers do not work well on multi-variable long sequence time series prediction problems[54]. This may be because multi-variable features will interfere Extraction of long-term sequence features[41]. Human activity perception based on wearable devices is essentially a long-sequence multi-variable time series analysis problem. Sensor data includes acceleration, gyroscopes, magnetometers, etc. channel signal. Inspired by[41], instead of using the classic transformer architecture, we introduce a channel-independent transformer architecture as the encoder-decoder of P2LHAP, each patch only contains data from a single wearable sensor channel information, each sensor signal sequences within the dataset undergoes processing independently within the Transformer model, allowing for the learning of diverse attention patterns tailored to each sequence. In contrast to the conventional architecture where all sequences are subject to the same attention mechanism, this individualized approach prevents potential interference caused by disparate behaviors within multi-channel sensor data. Moreover, by adopting a channel-independent strategy, the adverse impact of noise prevalent in certain sensor signal sequences can be mitigated as it no longer gets projected onto other sequences within the embedding space.
we utilize the original sensor data sequence along with the patch features acquired by the encoder to compute the representation of individual patches using the attention network for the prediction of activity labels assigned to each patch. Simultaneously, due to the variable size of the patch, the ultimate active label sequence may suffer from over-segmentation. To counter this issue, we devised a smoothing algorithm for the final Patch-level active label sequence, aimed at mitigating the over-segmentation phenomenon. This algorithm considers the number of activity categories surrounding the patch to adjust the current patch-level activity label.
The main contributions of this paper can be summarized as follows:
  • The concept of transforming patch sequences into label sequences was introduced for conducting activity perception tasks, where activity recognition, segmentation, and forecast are performed on the resulting activity label sequence. This approach has demonstrated enhanced performance across all three tasks.
  • Introduced a channel-independent Transformer architecture that can transform the Patch level sensor sequence at the current moment into an active label sequence corresponding to both the current and future Patch blocks.
  • Designed a smoothing method to analyze the number distribution of category labels surrounding each active label. This information is used to update the current active label, effectively reducing the occurrence of over-segmentation errors and providing smoother predictions within the same activity category.
  • The proposed P2LHAP method has been evaluated through experiments conducted on three widely used benchmark datasets for activity recognition, segmentation, and forecast. The experimental results demonstrate that our approach surpasses existing methods in terms of performance.
II Related Works
This section provides a comprehensive review of the progress made in related research on multi-task activity recognition, segmentation, and forecast utilizing sensor signals.
II-A Sensor-based activity recognition and segmentation tasks
Sensor-based human activity recognition (HAR) research primarily relies on inertial measurement unit (IMU) sensors found in wearable devices such as smart watches, smart insoles, and sports straps. The common approach for sensor-based activity recognition involves utilizing fixed-size sliding windows, which has shown impressive results[8]. [45] a long short-term memory (LSTM)-based classification method is proposed to identify fine-grained patterns according to the high-level features contextually extracted from the sequential motion data. Alternatively, [32] formulates HAR as a set forecast problem, where the prediction consists of ongoing active elements with varying cardinality. In the realm of HAR, [33] introduces an attention-based multi-head model for human activity recognition. Likewise, [46] propose a semi-supervised generative adversarial network (GAN) for CSI-based activity recognition (CsiGAN). Furthermore, [35] presents a triplet cross-dimensional attention method for sensor-based activity recognition tasks, deploying three attention branches to capture cross-interactions among sensor dimensions, time dimensions, and channel dimensions. Similarly, [36] introduces a novel multi-branch CNN that incorporates an attention mechanism to perform kernel selection among multiple branches with different receptive fields in the HAR domain. Additionally, [10] suggests a GAN-based approach to synthesize sensor data for sensor-based HAR tasks. In the context of realistic settings, [11] and [12] utilize sample evaluations to compare the performance of their recognition models. However, the length of the sliding window directly impacts the quality of feature extraction and classification accuracy [8]. Consequently, researchers have started focusing on adaptive sliding windows and have proposed sensor data segmentation strategies based on sliding windows. For instance, [13] introduces a novel multi-window fusion method that involves segmenting data using five different window sizes and fusing the recognition results in two stages. In a different approach, [14] presents a unified weakly supervised framework based on distributed kernel embeddings. This framework segments sensor streams, extracts powerful features from each segment, and trains a final classifier for activity recognition. Additionally, [15] suggests using a multivariate Gaussian distribution to detect signal differences between time windows. Nonetheless, defining the optimal window size remains a challenge due to the variation in activity durations.
Another effective method for sensor-based activity recognition is dense label prediction [4]. Inspired by the image semantic segmentation task, [18, 19, 16, 17] apply FCN (fully convolutional network) and U-net to sensor-based activity recognition to predict dense labels, achieving good recognition effects. [20] further proposes Conditional-UNet for coherent human activity recognition with multiple dense labels, explicitly modeling the conditional dependencies between these labels. Lastly, [33] explores the exploitation of potential relationships among multi-channel sensor modes and specific activities, investigating the effectiveness of data-agnostic enhancement of multi-modal sensor data streams for activity recognition prediction.
II-B forecast of activities using wearable sensors
Sensor-based human activity recognition systems have gained popularity with the widespread use of wearable devices in daily life. However, there is a notable lack of research focusing on sensor-based activity forecast. Recent studies have introduced attention models to predict multivariate motion signals from IMU sensors, thereby preserving pattern and feature information across multiple channels. Predictive models for these time series data can be utilized for human activity forecast by generating and classifying future activity signals. In [21], a combination of a neural network and Fourier transform was employed to convert the sensor activity data into an image format for predicting future signals. The system achieved a mean square error (MSE) value of 0.134 and was successfully used to predict elderly falls. In [22] introduced the SynSigGAN model, which utilizes bidirectional grid long short-term memory as the generator network of the GAN model, and a convolutional neural network as the discriminator network to generate future signals. Furthermore, recent work [23] proposes a Transformer-based GAN that can successfully generate real synthetic time series data sequences of arbitrary length, similar to real data sequences. Both the generator and discriminator networks of the GAN model are built using a pure Transformer encoder architecture, which outperforms traditional GAN structures based on RNN or CNN models. An additional study [41] presented the PatchTST time series forecast model, which is based on channel independence and patching, and demonstrated promising results. Although progress has been made in multivariate sensor time series data forecast, research on multi-category forecast of human activities based on multi-sensor signal channels remains limited.
II-C Sensor-based multi-task learning framework
Sensor-based multi-task activity recognition presents a significant challenge in the field. [24] employs recurrent neural networks to segment and identify activities and cycles using inertial sensor data. [25] proposes an efficient multi-task learning framework that utilizes commercial smart insoles. This framework addresses three health management-related problems by converting smart insole sensor data into recursive graphs and employing an improved MobileNetV2 as the backbone network. Experimental results demonstrate that this multi-task learning framework outperforms single-task models in tasks such as activity classification, speed estimation, and weight estimation. [26] introduces a multi-task model based on LSTM for predicting each activity and estimating activity intensity using wearable sensor signals. [6] combines the segmentation and classification tasks by passing the sensor signal through the MS-TCN module to predict sample-level labels and activity boundaries. This joint segmentation and classification approach calculates sample-level loss and boundary consistency loss to achieve improved performance. [5] proposes a novel multi-task framework called MTHARS, which leverages the dynamic characteristics of activity length to segment activities of varying durations through multi-scale window splicing. This approach effectively combines activity segmentation and recognition, leading to improved performance in both tasks.
III Method
Figure 2: Overview of the Proposed P2LHAP framework
In this section, we introduce our proposed method, which is a multi-task model for human activity recognition utilizing the Patch-to-Label Seq2Seq approach. Fig. 2 provides an overview of the model’s architecture. This model enables activity recognition, segmentation, and forecast tasks on continuous sensor data streams collected from wearable devices. Our approach consists of three primary components: 1) Patch-to-Label Seq2Seq framework, 2) Patching and channel independent Transformer architecture, and 3) activity classification, segmentation, and forecast module.
III-A Patch-to-Label Seq2Seq framework
[…]
III-B Patching and channel independent Transformer architecture
III-B1 Patching and channel independent method
We process each sensor signal channel independently, specifically, We segment the input sensor signal sequence $\mathbf{X}_{1:L}$ into $M$ channel sensor signal sequences $\mathbf{x}^{(i)}$ based on its dimension $M$, $\mathbf{x}^{(i)}$ represents the sensor sequence of the $i$-th channel, $i\in\{1,2,\cdots,M\}$. we normalize each $\mathbf{x}^{(i)}$ before patching and the mean and deviation. Next, we divide $\mathbf{x}^{(i)}$ into overlapping or non-overlapping Patches and the length of the Patch block is $P$. The overlapping part is its step size, represented by $S$. By segmentation, a patch-level sensor sequence $\mathbf{x}_{p}^{(i)}\in\mathbb{R}^{1\times N}$ will be obtained, where $N=\left\lfloor\frac{(L-P)}{S}\right\rfloor+2$. $\mathbf{x}_{p}^{(N)}$ is the last patch of the segmented patch sequence, this patch is learned through the previous $N-1$ patches, a sequence representation into the future is obtained, with its initial value filled by repeatedly replicating the last number of the original sequence. Finally, we add the corresponding position coding to $\mathbf{x}_{p}^{(i)}$ to obtain $\mathbf{x}_{d}^{(i)}$, and then transmit the time series of each channel to the Transformer backbone respectively.
III-B2 Encoder
We use the native Transformer encoder [dosovitskiy2020image] to model the features of patch-level sensor sequences, similar to the method used in [41]. The Encoder module utilizes trainable linear projections to map continuous Patch blocks to a D-dimensional Transformer latent space $\mathbf{W}_{p}\in\mathbb{R}^{D\times P}$. It incorporates a learnable additive positional encoding $\mathbf{W}_{pos}\in\mathbb{R}^{D\times N}$ to capture the temporal order of the patches, represented by $\mathbf{x}_{d}^{(i)}=\mathbf{W}_{p}\mathbf{x}_{p}^{(i)}+\mathbf{W}_{pos}$, where $\mathbf{x}_{d}^{(i)}\in\mathbb{R}^{D\times N}$ represents the input fed into the Transformer encoder. For each head $h=1,\cdots,H$ in the multi-head attention, the Patch blocks are transformed into query matrices $\mathbf{Q}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{Q}$, key matrices $\mathbf{K}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{K}$, and value matrices $\mathbf{V}_{h}^{(i)}=(\mathbf{x}_{d}^{(i)})^{T}\mathbf{W}_{h}^{V}$, where $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K}\in\mathbb{R}^{D\times d_{k}},\mathbf{W}_{h}^{V}\in\mathbb{R}^{D\times D}$. Subsequently, the attention outputs are normalized to obtain the final attention output $\mathbf{Q}_{h}^{(i)}\in\mathbb{R}^{D\times N}:$
$(\mathbf{Q}_{h}^{(i)})^{T}=Attention(\mathbf{Q}_{h}^{(i)},\mathbf{K}_{h}^{(i)},\mathbf{V}_{h}^{(i)})$ (1)
$Attention(\mathbf{Q},\mathbf{K},\mathbf{V})=softmax(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})\mathbf{V}$ (2)
The multi-head attention module also includes a BatchNorm layer and a feed-forward network with residual connections.
III-B3 Decoder
The Decoder module combines the representation $\mathbf{Z}^{(i)}=\left\{\mathbf{z}_{1},\cdots,\mathbf{z}_{p}\right\}$, derived from the encoder output, with the corresponding Patch-level wearable sensor data $\mathbf{X}^{{}^{\prime}}=\left\{\mathbf{x}_{1}^{{}^{\prime}},\cdots,\mathbf{x}_{p}^{{}^{\prime}}\right\}$ to generate the activity label sequence using a key-query-value attention mechanism. By multiplying the weight matrices $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ with the Patch-level wearable sensor data, we obtain representations of the key and value, denoted as $\mathbf{K}=\mathbf{W}_{k}\mathbf{X}^{{}^{\prime}}$ and $\mathbf{V}=\mathbf{W}_{v}\mathbf{X}^{{}^{\prime}}$, respectively. At the same time, the query vector $\mathbf{Q}$ is obtained by multiplying the feature representation $\mathbf{z}^{(i)}$ from the encoder with the weight matrix $\mathbf{W}_{q}$, expressed as $\mathbf{Q}=\mathbf{W}_{q}\mathbf{Z}^{(i)}$. Subsequently, the inner product of $\mathbf{Q}$ and $\mathbf{K}$ is passed through a temperature softmax layer to obtain attention scores, which are then used to weigh the summation of the value vector $\mathbf{V}$, resulting in the representation of $\mathbf{z}^{{}^{\prime}(i)}:$
$\mathbf{z}^{{}^{\prime}(i)}=softmax(\mathbf{Q}^{T}\mathbf{K}/\tau)\mathbf{V}=f_{\theta}(\mathbf{X}^{{}^{\prime}},\mathbf{Z}^{(i)})$ (3)
Here $\tau$ is the temperature parameter, and $f_{\theta}(\cdot)$ represents the attention mechanism network.
III-C Activity classification
we employ a Flatten layer with a linear head and a Softmax layer to generate the final human activity label sequence $\hat{y}$. The representation of $\hat{y}$ is summarized as follows:
$softmax(x)=\frac{e^{x_{i}}}{{\textstyle\sum_{j=1}^{n}e^{x_{j}}}}$ (4)
$\hat{y}=argmax_{i}(softmax(Linear(\mathbf{z}^{{}^{\prime}}))$ (5)
During the training phase, we utilize patch-level label predictions and true labels to construct a patch-level classification loss, aiming to effectively optimize the model parameters. This loss is defined as follows:
$L_{cls}=l_{p}(\hat{y},y)$ (6)
Among them, $\hat{y}$ and $y$ represent the patch-level prediction label and the real label, respectively. Given the imbalance of human activity classes under unconstrained conditions, we employ the Softmax CrossEntropy loss function $l_{p}(\cdot)$ and consider the number of effective samples for classification. The calculation formula is as follows:
$l_{p}(\hat{y},y)=-\frac{1}{L}\sum_{l=1}^{L}\sum_{c=1}^{C}ylog(\hat{y})$ (7)
Where $L$ signifies the length of input sequence, and C represents the activity class number of dataset. In sensor-based HAR tasks, we leverage this loss function and consider the number of effective samples to amplify the weight assigned to minority classes. This approach aims to yield fairer and more robust results for each individual class.
III-D Activity segmentation and smoothing strategy
Given that the classification loss treats each patch independently, it may result in unwarranted over-segmentation effects. To achieve smoother transitions between patches and minimize over-segmentation errors, we employ the truncated mean square error as our chosen smoothing loss function. The formulation of the smoothing loss is as follows:
$L_{seg}=L_{T-MSE}=\frac{1}{T_{c}C}\sum_{t,c}\widetilde{\Delta}_{t,c}^{2}$ (8)
$\widetilde{\Delta}_{t,c}^{2}=\left\{\begin{matrix}\Delta_{t,c}\quad,\Delta_{t,c}\leq\tau\\ \quad\tau\quad,otherwise\end{matrix}\right.$ (9)
$\Delta_{t,c}=\left|log\widetilde{y}_{t,c}-log\widetilde{y}_{t-1,c}\right|$ (10)
$C$ represents the activity category, $T_{c}$ denotes the number of samples belonging to the c-th category, and $\widetilde{y}_{t,c}$ signifies the probability of the activity category c being represented by the patch at time t and $\tau$ is the smoothing threshold($\tau=2$).
For the activity label sequence $\hat{y}$ output by the model, in order to reduce the impact of over-segmentation on subsequent activity merging, we set each activity label as the center, set a smoothing window with a size of smooth_size, and set the starting position to the smoothing window Half of the activity label sequence within the window, we count and sort the activities of each category, and select the activity category with the most surrounding distribution as the current activity label according to the sorting result. After updating the activity label sequence, the performance of activity recognition and segmentation. Finally, it is enhanced by linking Patches of the same activity class to determine the start and end position of each activity in the data flow. We set the boundary set of each activity $B=\{(s_{i},e_{i})\}_{i=1}^{N}$, where $s_{i}$ represents the start subscript of the i-th activity, $e_{i}$ represents the end subscript of the i-th activity, and N is the number of activities in the activity label sequence. The activity set corresponding to each boundary is represented by $\hat{a}=\{a_{i}\}_{i=1}^{N}$, where $a_{i}$ is the activity corresponding to the i-th boundary. The Process of Algorithm is described in the Algorithm 1.
III-E Activity forecast
Within the activity forecast module, the continuous input Patch blocks retain one Patch block at the end. This involves filling P repeated values of the last $\mathbf{x}_{i}\in\mathbb{R}$ value at the end of the original sequence. The resulting sequence of generated Patches, $\mathbf{x}_{p}^{(i)}$, is then represented as $\mathbf{z}^{{}^{\prime}(i)}\in\mathbb{R}^{D\times N}$ through the Transformer backbone. By utilizing a linear layer and a fully connected layer with Softmax as the activation function, we obtain the predicted activity sequence, $\hat{y}_{pre}={\hat{y}_{L+1},\cdots,\hat{y}_{L+T_{P}}}\in\mathbb{R}^{1\times T_{P}}$, using the representation of the last patch. It is important to note that $T_{P}$, the length of the future segment, is determined by the number of Patch blocks within the time $T_{pre}$, specifically $T_{p}=\frac{T_{pre}}{P}$.
Algorithm 1 Segmentation Algorithm
0:  Sensor dataset: $\bf{X}$
0:  Activity label sequence segmentation result: $\hat{y}$       Activity boundaries: $B=\{(s_{i},e_{i})\}_{i=1}^{N}$       Activity class: $\hat{a}=\{a_{i}\}_{i=1}^{N}$
1:  $\hat{y}\leftarrow Model(\bf{X})$ by Eq.(1), (2), (3), (4) and Eq(5)
2:  $smooth\_size$leftarrow smoothing window size
3:  set a smooth sequence of active label: smooth__\__data $\leftarrow$ []
4:  $half\_size\leftarrow smooth\_size//2$
5:  for $i=0$ to $len(\hat{y})$  do
6:     start $\leftarrow$ max(0, $i-half\_size$)
7:     end $\leftarrow$ min(len($\hat{y}$, $i+half\_size+1)$)
[…]
9:     add new_activity_label to smooth__\__data
10:  end for
11:  $\hat{y}$leftarrow smooth__\__data
12:  current activity type a $\leftarrow$ $\hat{y}[0]$
13:  initial activity set index $j$
14:  activity start and end index s, e $\leftarrow$ 0
15:  for $i$ $=$ $1$ to $len(\hat{y})$  do
16:     if $\hat{y}_{i}$neqthen
17:        Add $a$ to $\hat{a}_{j}$
18:        $a$leftarrow $\hat{y}_{i}$
19:        $e$leftarrow $i$
20:        Add ($s$, $e$) to $B$
21:        $s$leftarrow $e$ + 1
22:        $j$leftarrow $j$ + 1
23:     end if
24:  end for
In the training stage, based on the predicted activity label sequence $\hat{y}_{pre}$ and the ground-truth activity sequence $y_{gt}$, all patch-level prediction losses can be expressed as:
$L_{pre}=l_{p}(\hat{y}_{pre},y_{gt})$ (11)
Finally, the previously mentioned loss functions from Equations (6), (8), and (11) are incorporated into the equation, resulting in the final loss function:
$L=L_{cls}+L_{seg}+L_{pre}$ (12)
We have introduced the main structure of P2LHAP. Algorithm 2 summarizes the training process of P2LHAP.
IV EXPERIMENTS AND RESULTS
In this section, we provide a comprehensive overview of our experiments and utilize cutting-edge research to evaluate our proposed multi-task framework. Firstly, we present the experimental settings, datasets, and evaluation metrics employed in our study. Subsequently, we perform a series of experiments to assess the effectiveness of our method, including qualitative and quantitative overall performance analysis as well as ablation experiments conducted using the Patch strategy.
Algorithm 2 The training process of P2LHAP
0:  Training dataset: $\{\textbf{X}_{train},y_{train}\}$      Validation dataset: $\{\textbf{X}_{val},y_{val}\}$     Test dataset: $\{\textbf{X}_{test},y_{test}\}$
0:  Patch-level activity label predictions: $\hat{y}$
[…]
2:  while $e\leq Epoch_{max}$ and $Convergent$ do
3:     Get a mini-batch data $\textbf{X}_{batch}$ in $\textbf{X}_{train}$
4:     $\hat{y}\leftarrow Model(\textbf{X}_{batch})$ by Eq.(1), (2), (3), (4) and Eq.(5)
5:     $y_{gt}\leftarrow y_{train}$
6:     $\hat{y}_{pre}$leftarrow $\hat{y}[-T_{p}:-1]$
7:     Calculate $L_{cls}$ with eq.(6), (7)
8:     Calculate $L_{pre}$ with eq.(11)
9:     Get $\hat{y}_{seg}$ through algorithm 1
10:     $\hat{y}$leftarrow $\hat{y}_{seg}$
11:     Calculate $L_{\text{seg}}(\hat{y}$, $y_{\text{gt}}$) with eq.(8)
12:     e $\leftarrow$ e + 1
13:     Update model parameters based on eq.(12)
14:     $adjust\_learning\_rate(optimizer,e,args)$
15:     Calculate $\hat{y}_{val}$ with Eq.(6), (8) and check convergence
16:  end while
17:  Update $\hat{y}$ from test sequence $\textbf{X}_{test}$ based on the model
IV-A Experimental setup
Our method was implemented using PyTorch and trained on an RTX 3090 GPU. The batch size was set to 64, and we utilized the Adam optimizer with a learning rate of 0.0001.
For classification, segmentation, and forecast tasks, we employed the entire continuous dataset as input and divided it into training, validation, and test sets with a ratio of 0.7, 0.1, and 0.2 respectively. To evaluate the classification and segmentation performance of the model, we compared the forecast output with the real label. In the forecast task, the forecast patch length was set to 8, and performance assessment was based on a comparison between the forecast output patch label and the real label.
1) Evaluation metrics: For the evaluation of the obtained Patch-level label sequences, we use the following metrics:
1. Activity classification
Accuracy: Accuracy is the overall measure of correctness across all classes and is calculated as follows:
$Accuracy=\frac{\sum_{c=1}^{C}TP_{c}+\sum_{c=1}^{C}TN_{c}}{\sum_{c=1}^{C}(TP_{c}+TN_{c}+FP_{c}+FN_{c})}$ (13)
Weighted F1 score: The weighted F1 score is calculated based on the proportion of Patch samples:
$F_{w}=\sum_{c=1}^{C}2*w_{c}\frac{prec_{c}\cdot recall_{c}}{prec_{c}+recall_{c}}$ (14)
Where $prec_{c}=\frac{TP_{c}}{TP_{c}+FP_{c}}$ and $recall_{c}=\frac{TP_{c}}{TP_{c}+FN_{c}}$ represent precision and recall, respectively. The weight $w_{c}$ is calculated as the ratio of $n_{c}$ to $N$, where $n_{c}$ is the number of Patch samples per class, and $N$ is the total number of Patch samples.
2. Activity segmentation
Jaccard Index: The proposed metric assesses the degree of overlap between actual activity segments and predicted activity segments, specifically focusing on the sensitivity towards over-segmentation errors. For the purpose of representation, $G_{i}$ and $P_{i}$ are introduced as respective variables denoting the actual and predicted activity ranges. In this context, the subscript i signifies the ith segment within class C. The calculation of this metric can be determined utilizing the prescribed formula:
$JI=\frac{1}{C}\sum_{c=1}^{C}\frac{G_{i}\cap P_{i}}{G_{i}\cup P_{i}}$ (15)
3. Activity forecast
Mean Square Error(MSE): MSE serves as a prevalent metric to evaluate the performance of forecast models. It computes the average of the squared differences between predicted values and actual values, thereby providing a quantification of the degree of deviation between the predicted and actual values. A lower MSE signifies a more precise alignment between the model’s predictions and the actual values, indicating superior performance. The formula for calculating MSE can be represented as follows:
$MSE=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (16)
IV-B Datasets
To evaluate our proposed method, we adopt three challenging public HAR datasets: WISDM, PAMAP2, UNIMIB SHAR. A brief introduction to the dataset follows.
WISDM Dataset[27]: The dataset was collected from 29 participants using mobile phones equipped with triaxial accelerometers placed in their pants pockets. The data was sampled at a frequency of 20 Hz. Participants engaged in six activities on a daily basis: walking, strolling, ascending stairs, descending stairs, standing still, and standing. To address missing values in the dataset, the mean of the corresponding column was used for filling them.
PAMAP2 Dataset[28]: The dataset consists of data collected from nine participants who wore IMUs on their chest, hand, and ankle. These IMUs collected information on acceleration, angular velocity, and magnetic sensors. Each participant was instructed to complete 12 mandatory activities such as lying down, standing, going up and down stairs, as well as 6 optional activities like watching TV, driving, and playing ball.
UNIMIB SHAR Dataset[29]: The dataset was collected by researchers at the University of Milan-Bicocca. Thirty volunteers, ranging in age from 18 to 60 years old, were equipped with Bosch BMA220 sensors on their Samsung mobile phones. The phones were placed in both the participants’ left and right pockets, resulting in a total of 11,771 activities recorded. The dataset comprises 17 categories of activities, categorized into two coarse-grained groups: 9 activities of daily living (ADL) and 8 falling behaviors.
IV-C Comparison algorithm
We compare our method with state-of-the-art algorithms for activity recognition, segmentation, and forecast.
  • MPCC-MRCSE[58]: A multidimensional parallel convolutional connected (MPCC) deep learning network based on MMS data for HAR is proposed that fully utilizes the advantages of multidimensional convolutional kernels.
  • CNN-BiLSTM[59]: Proposed multibranch CNN-BiLSTM network does automatic feature extraction from the raw sensor data with minimal data pre-processing
  • A&D[11]: A deep learning method for HAR that includes modules such as CIE, AGE, and center loss. The method achieves state-of-the-art recognition performance through data enhancement techniques.
  • MRFC-VAE-Net [56]: A multi-resolution fusion convolutional network is proposed to make up for the shortcoming that confusing activities are difficult to correctly identify, thereby improving the accuracy of identification.
  • Multi-ResAtt[60]: Multi-ResAtt integrates a recurrent neural network with attention to extract time-series features and perform activity recognition.
  • ConvAttention[33]: A lightweight multi-head model that incorporates an attention mechanism to enhance the representation capability of CNNs for human activity recognition.
  • CE-HAR[57]: Provide an alternative called Channel Equalization to reactivate these inhibited channels by performing whitening or decorrelation operation, which compels all channels to contribute more or less to feature representation
  • TripleCrossAttention[35]: This method introduces a triple cross-dimensional attention mechanism for sensor-based activity recognition tasks.
  • SK[36]: A multi-branch CNN model that utilizes selective kernel mechanisms for HAR. Additionally, it incorporates attention mechanisms for the first time in the HAR domain to select kernels between multiple branches with different receptive fields (RFs).
  • MTHARS[5]: A multi-task deep neural network that enhances activity segmentation and recognition by utilizing a multi-scale window splicing method. This approach effectively combines both tasks and improves performance.
  • BCMLF-HARS[6]: A multi-task learning framework for joint segmentation and recognition of activities using wearable sensors. It introduces a multi-level temporal convolutional network (MS-TCN) to predict activities at the sample level, overcoming the challenge of multi-class windows. The consistency loss attention mechanism effectively enhances the consistency between activity and boundary predictions.
IV-D Comparisons with activity recognition algorithm
In this section, we compare our method with other competing methods that have been evaluated on three publicly available datasets for wearable-based Human Activity Recognition (HAR). To conduct a thorough performance analysis, we use non-overlapping patches of size 10 to compare the overall recognition performance.
We compared the recognition performance of our proposed method with other similar algorithms as mentioned earlier. The recognition performance of various competing methods is summarized in Table I. Our method demonstrated improvements in recognition accuracy across three publicly available HAR datasets. Particularly, on the Unimib and PAMAP2 datasets, our method achieved the highest recognition performance, obtaining weighted average F1 scores of 82.04% and 98.92% respectively. Remarkably, our method surpassed the performance of the leading competing method by 3.26% in terms of weighted average F1 score on the Unimib dataset.
[…]
Methods Datasets PAMAP2 WISDM UNIMIB SHAR
Acc F1 Acc F1 Acc F1
P2LHAP 98.71% 98.92% 97.51% 97.52% 81.51% 82.04%
Other Reachers 90.70% 90.55% [11] 92.71% 92.63% [34] 76.30% 75.38% [35]
93.80% 93.38% [36] 96.04% 96.05% [59] 75.89% 74.63% [36]
94.27% 94.25% [59] 96.85% 97.20% [33] 76.48% 75.71% [5]
93.19% 92.96% [60] 97.51% 97.25% [36] 77.29% 76.83% [60]
94.50% 94.80% [5] 99.04% 99.18% [57] 78.65% 78.78% [57]
92.14% 92.18% [57] 97.11% 98.64% [56]
94.59% 94.47% [6]
97.39% 98.12% [56]
- 98.33% [58]
Figure 3: The confusion matrices on the PAMAP2 dataset for patch sizes of 10 and 200 (a) $Patch\_size$=10, (b) $Patch\_size$=200
In Fig. 3, we present the confusion matrix of our method for activity classification on the PAMAP2 dataset using patch sizes of 200 and 10, respectively. The confusion matrix provides insight into the accuracy of our method for various activity categories. It is evident that, despite the different patch sizes, there is minimal variation in overall accuracy. However, when the patch size is small, over-segmentation occurs, necessitating active segmentation through dividing the patch blocks. This approach enhances the recognition performance for most activity categories.
Figure 4: Segment visualization results on PAMAP2 and WISDM datasets
IV-E Qualitative segmentation results
To demonstrate the segmentation performance of our proposed method, this section presents a visualization of patch-level activity predictions in Fig. 4, with each color representing a different activity category. We compared sequence fragments obtained through model segmentation with selected sequence fragments from the PAMAP2 and WISDM datasets. In practical application scenarios, excessive segmentation can adversely affect prediction accuracy. Nonetheless, as depicted in the figure, our method successfully reduces most over-segmentation errors, leading to more precise and consistent patch-level predicted segments within each active segment.
Table II displays the segmentation performance of our proposed method in comparison to competing algorithms. Our method excels in both segmentation and recognition tasks, consistently outperforming the comparison algorithms. Specifically, our Jaccard Index results on the Pamap2 dataset surpassed those of A&D and BCMLF-HARS by 9.82% and 3.47% respectively, Where (w/o) represents the segmentation Jaccard Index scores obtained without using the smoothing strategy.
TABLE II: RESULTS OF SEGMENTATION METRICS ON PAMAP2 DATASET.
Seg Metric PAMAP2 DATASET
Attn A&D BCMLF-HARS ours(w/o) Ours
Jaccard Index 0.7908 0.8312 0.8947 0.8445 0.9294
IV-F Patch prediction analysis
In this section, we assess the performance of the activity prediction module using our model to forecast the duration of activity categories in the next 8 patch blocks. Fig. 5 displays the confusion matrix for the future activity prediction in the pamap2 dataset. Fig. 5 (a),(b),(c) present the prediction confusion matrices of Conv2LSTM[38], Seq2Seq-LSTM[39], and Seq2Seq-LSTM-PE-MA[40] predictors for five activity categories in the dataset, while Fig. (d) illustrates the predictions of our proposed model. The confusion matrix results for all campaigns indicate an average accuracy of 96.59%. Additionally, we compare the predicted outputs of future sensor signals with the ground truth signals by evaluating their mean squared error (MSE). To enable the model to predict future sensor signals, we exclude the last classification header and utilize the first half of the input sensor sequence for model input, reserving the second half to match the size of the predicted output data. Table III presents the performance comparison of our model with five predictors for sensor sequence signal prediction.
Figure 5: Confusion matrix of our model and three other models in predicting activity classification results on the PAMAP2 dataset
TABLE III: Evaluation results of the forecasters.
Metric Models
SynSigGAN Conv2LSTM Seq2Seq-LSTM TTS-GAN Ours
MSE 0.355 0.302 0.240 0.163 0.126
IV-G Ablation experiments
In this section, we conduct two ablation experiments to investigate the influence of varying patch block sizes on model performance and the efficacy of patching operations in generating label sequences.
Impact of Patch Size: various sizes and quantities of patches can yield different outcomes. Hence, we employ a scaling approach with multiple patches to analyze their performance. We provide the comprehensive results in Table IV. We observed a decrease in accuracy when the Patch size is reduced significantly. This decline can be attributed to the issue of excessive segmentation. However, when the Patch size is set to 10, it consistently exhibits exceptional recognition performance across all three datasets.
The impact of Patching modules: we examined the effects on the results by not patching the sensor data stream, but instead using the Transformer backbone directly. The F1 score obtained on the benchmark dataset is presented in Table V. It was observed that when the data was not patched during processing, the accuracy of the experimental results deteriorated. Hence, applying data patching enhanced the performance of activity recognition.
TABLE IV: $F_{1}$ values with different patch sizes on benchmark datasets.
Datasets Patch Size
P=1 P=2 P=5 P=10 P=20
   WISDM
0.8544 0.9056 0.8819 0.9751 0.9568
   PAMAP2
0.9151 0.9406 0.9610 0.9892 0.9667
   UNIMIB
0.6719 69.50 75.66 0.8204 74.76
[…]
Model WISDM PAMAP2 UNIMIB
only transformer backbone 0.7090 0.8783 0.7533
P2LHAP $\bf{0.9752}$ $\bf{0.9892}$ $\bf{0.8204}$
V CONCLUSION AND FUTURE WORK
In this paper, we propose P2LHAP, an activity-perception model that offers a patch sequence to label sequence model design solution tailored to identify, segment, and forecast human activities through sensor data analysis. Unlike conventional Transformer-based models, P2LHAP partitions the native sensor data stream into multiple patch blocks and utilizes a channel-independent approach to mitigate noise interference among sensor channels. The forecast of forthcoming activity sequences is achieved through reasoning with a Transformer encoder-decoder. To address the issue of over-segmentation, we refine labels based on the nearby distribution of labels close to the active label, thus reducing clutter in the distribution of active labels. Leveraging the P2LHAP framework, we effectively tackle activity segmentation, recognition, and forecast within the realm of activity perception tasks. Our experimental evaluation on three standard datasets demonstrates that our method surpasses existing approaches in classification, segmentation, and forecast tasks.
In future work, we will explore extending this model to address cross-human domain challenges, using self-supervised methods to reduce the workload of activity labeling. Additionally, to mitigate the quadratic time complexity issue stemming from the Transformer architecture, we aim to investigate employing a lighter backbone network architecture for undertaking activity-perception tasks like activity recognition, segmentation, and forecast. This approach will allow our architecture to be deployed on mobile devices for real-time applications.